OpenStudy (anonymous):

What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point: (8, -8); slope: 3.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@DDS

jimthompson5910 (jim_thompson5910):

point slope form is y-y1 = m(x-x1)

jimthompson5910 (jim_thompson5910):

m is the slope (x1,y1) is the given point

OpenStudy (anonymous):

IS it y- 8 =3 (x+8)?

jimthompson5910 (jim_thompson5910):

notice how it's y-y1 if y1 is -8, then y-y1 y-(-8) y+8 see how this works?

OpenStudy (anonymous):

Yeah. So, am I wrong or right?

jimthompson5910 (jim_thompson5910):

look above and you tell me

OpenStudy (anonymous):

right?

jimthompson5910 (jim_thompson5910):

you said "IS it y- 8 =3 (x+8)" but I said "notice how it's y-y1..." (see the post above) you'll see the difference

OpenStudy (anonymous):

can you just tell me if it's right or wrong..? -_-

jimthompson5910 (jim_thompson5910):

it's wrong, do you see how?

OpenStudy (anonymous):

no.

jimthompson5910 (jim_thompson5910):

look at how I got y+8 you got y-8

OpenStudy (anonymous):

so, thats the only difference?

jimthompson5910 (jim_thompson5910):

no there is one more

jimthompson5910 (jim_thompson5910):

it has to do with the x-x1 term

OpenStudy (anonymous):

IT WOULD BE X-8?

jimthompson5910 (jim_thompson5910):

better

OpenStudy (anonymous):

Is it y+8=3(x-8)?

jimthompson5910 (jim_thompson5910):

yes it is